OpenStudy (anonymous):

Maldo Geronimo conducts two-day training sessions for an oxygen equipment company. When he traveled last month, his lodging was $235 per night for two nights and breakfast and lunch were $46 per day. His total transportation expenses were $535. He has a budget of $1,200. What is the maximum he can spend on dinner each of the two nights? Express your answer in dollars and cents.

OpenStudy (anonymous):

Hey there! Actually, I just figured out the answer to this, and have already check that it is correct. Here's what I did: (235*2)+(46*2)+535=1,097 1200-1097=103 103/2=51.5 The answer would then be $51.50